TRANSIENT FREQUENCY-MODULATION DIODE LASER SPECTROSCOPY OF METHYLENE

\underline{BOR-CHEN CHANG}, Department of Chemistry, National Central University, Chung-Li, Taiwan 32054; ANDREW J. MARR and TREVOR J. SEARS, Department of Chemistry, Brookhaven National Laboratory, Upton, New York 11973-5000.

Transient frequency-modulation (FM) techniques have been combined with diode laser absorption to acquire high resolution spectra of the methylene (CH_{2}) b̃^{1}B_{1} ←ã^{1}A_{1} electronic transition in the vicinity of 10000cm^{−1}-10600cm^{−1}. These spectra represent the first application of transient FM techniques using a simple and inexpensive diode laser source to high resolution spectroscopy. By comparison with previous data obtained using a dual-beam absorption setup, the improvement in signal-to-noise ratio is dramatic. A sensitivity (ΔI/I) for raw signals of 1.1 × 10^{−4} in 0.6 μsec is now routinely achieved for 0.4 mW laser power at the detector. The integrated absorption signals have even larger signal-to-noise ratios. This can be compared with the shot noise limit for absorption of 2.9×10^{−5} for these parameters. This talk will concentrate on details of the experimental design. Acknowledgments: The experimental work was carried out at Brookhaven National Laboratory under Contract No. DE-AC02-76CH00016 with the U.S. Department of Energy and supported by its Division of Chemical Sciences, Office of Basic Energy Sciences. BCC also thanks the National Science Council, Taiwan for its support of this work under Grant No. NSC87-2113-M-008-